# Beyond Image Prior: Embedding Noise Prior into Conditional Denoising Transformer

### v1

@@ 3.3 Conditional Denoising Transformer
Due to the strong capability of Transformer-based denoising models (e.g., CAT [49], Uformer [8], Restormer [9] and etc), we design a Conditional denoising Transformer (Condformer) by embedding the noise prior into the self-attention module. Particularly, as the independence of noise prior and image prior, guiding the model to learn from image prior and noise prior separately is the primary principle in designing the Condformer.
@@ 3.3.1 Embedding noise prior in latent space
Existing image denoising networks typically use a global residual connection for noise prediction and an U-shape encoder-decoder structure for image prior representation, suggesting that noise is implicitly concealed in the latent space as illustrated in Fig. 5. In latent space manipulation [67, 68], the latent space code from the encoder effectively represents implicit semantic attributes for domain transfer. Inspired by this and considering the residual attribute of noise, the latent space code of a noisy image significantly represents noise statistics.
As mentioned earlier in Section 1, an effective conditional denoising model should separately consider the image and noise priors, adhering to the principle of their independence. Specifically as shown in Fig. 5, we firstly extract the latent space code $\mathbf{X}$ of the noisy image using a denoiser encoder, and then construct a feature fusion module to embed the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$.
@@ 3.3.2 Overall pipeline
Considering the efficiency in practical applications, we employ the Restormer as our denoiser baseline, replacing its self-attention module by a conditional self-attention module that embeds noise prior in the latent space. Following [9], given a noisy color observation $\mathbf{I}_{\text{rgb}}\in\mathbb{R}^{3\times h\times w}$, we construct a multi-scale hierarchical denoiser encoder which consists of three channel-wise Transformer blocks to capture cross-covariance across channels, generating the latent space code $\mathbf{X}\in\mathbb{R}^{8c\times\frac{h}{8}\times\frac{w}{8}}$, where $c$ is the number of channels.
@@
$\hat{\mathbf{X}}=\text{CondSA}(\mathbf{X},\mathbf{z})$ (16)
This is then fed into a feed-forward network (FFN) for feature transformation. In particular, to effectively exploit the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ estimated by LoNPE algorithm or network, CondSA encodes this prior into a latent conditional embedding vector $\mathbf{z}\in\mathbb{R}^{1\times k}$ using a shallow module with fully-connected layers. Each CondSA has a specific embedding vector to adapt the intermediate features. The rectified latent space code, equipped with noise prior, guides the denoiser decoder for specific denoising. In particularly, to preserve fine structural and textural details in the restored images, we use a hierarchical skip-connection strategy [9] that integrates low-level features from the encoder and high-level features from the decoder.
@@ 3.3.3 Conditional Self-Attention Mechanism
where $\mathbf{X}$ and $\mathbf{Y}$ are the input and output features. $\mathbf{Q}\in\mathbb{R}^{c\times hw}$, $\mathbf{K}\in\mathbb{R}^{c\times hw}$ and $\mathbf{V}\in\mathbb{R}^{c\times hw}$ indicate the query, key and value matrix obtained by encoding the input feature $\mathbf{X}\in\mathbb{R}^{c\times h\times w}$ with linear layers $W^{Q}$, $W^{K}$ and $W^{V}$ respectively, each of which stacks a $1\times 1$ convolution and a $3\times 3$ depth-wise convolution layer. Besides, $\alpha$ is a learnable scaling parameter to control the magnitude of the cross-covariance of $\mathbf{Q}$ and $\mathbf{K}$ before applying a softmax layer.
@@
where $\mathbf{z}\in\mathbb{R}^{1\times 2k}$ represents the conditional embedding vector from the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ by repeating $k$ times in channel dimension.
The conditional attention should effectively represent the correlation between-in the intermediate image features, and the latent correlation between the intermediate image features and noise prior. Therefore, a fusion module is necessary for correlation representation of noise and image priors. Intuitively, the query tensor $\mathbf{Q}$ and the key tensor $\mathbf{K}$ indicates the information for feature retrieval [42]; instead, the value tensor $\mathbf{V}$ represent the property of the input feature. Therefore, it is reasonable to embed noise prior into the query/key tensors and generate the conditional counterparts $\mathbf{Q}^{\prime}$ and $\mathbf{K}^{\prime}$, so we have
@@
where $\Phi_{\text{LFM}}$ represents a linear fusion module (LFM) layer for feature fusion. In detail, in order to fuse the query/key tensor and the conditional embedding vector in channel and spatial dimension, we can simply concatenate them in channel-wise and feed them into a a $1\times 1$ convolution and a $3\times 3$ depth-wise convolution layer as the convention. In this way, the noise prior would be embedded into each channel and a local receptive field of the tensors.
Figure 6: Noise prior estimation on synthetic noises with Poisson-Gaussian distribution. By applying our LoNPE algorithm, we can effectively calculate a relative accurate noise prior from a single noisy observation.
@@ 4 Experiments
We first describe the experimental setup and then evaluate the performances of our LoNPE and Condformer on noise statistics and blind image denoising. Finally, we perform ablation studies to demonstrate the effectiveness of our methods.
@@ 4.1.1 Datasets and Metrics
Synthetic Datasets. Following [53], we adopt several sRGB image datasets for training the LoNPE and Condformer networks, including the DIV2K and Flicker2K dataset [69] with 3650 high-quality 2K images, the Berkeley segmentation dataset (BSD) [70] with 400 images, Waterloo Exploration Database (WED) dataset [71] with 4744 images. Considering both signal-dependent and signal-independent noises, we randomly add noises on the clean sRGB image with Poisson-Gaussian noise model. The noise level are set to $\sigma_{\text{s}}\sim[0,0.3]$ and $\sigma_{\text{r}}\sim[0,50/255]$, both of which indicate the groundtruth noise prior to demonstrate the effectiveness of our LoNPE algorithm. Especially, the Poisson-Gaussian noise will be degraded to additive Gaussian white noise (AWGN) when $\sigma_{\text{s}}=0$. To evaluate the performance, we apply our methods on several benchmarks, including CBSD68 [70], Kodak24 [72] and Urban100 [73].
Real Datasets. Consistent with previous real image denoising work [9], we adopt the SIDD-Medium dataset [10] for real noise statistics and real image denoising tasks, which contains 320 noisy-clean image pairs in both raw and sRGB domains. Particularly, we first apply our LoNPE algorithm on 320 noisy raw images to estimate their corresponding noise priors, and then utilize them to help training the LoNPE and Condformer networks on the 320 sRGB noisy-clean image pairs. For validation, 1024 pairs of noisy-clean sRGB image patches from SIDD validation dataset [10] are adopted. Besides, evaluation is also conducted on 1280 noisy $256\times 256$ patches from the SIDD benchmark dataset [10] and 50 noisy $512\times 512$ images from the DND benchmark dataset [12].
Evaluation Metrics. Two commonly-used image quality assessment criteria are adopted to evaluate the performances: Peak Signal-to-Noise Ratio (PSNR) and Structural SIMilarity (SSIM) [74], and are calculated in the sRGB domain. Note that, since the clean groundtruthes of SIDD and DND benchmark datasets are unavailable, we calculate these metrics on only SIDD validation dataset and obtain the metrics of them from the online servers [^1] SIDD: https://abdokamel.github.io/sidd/;
@@ 4.1.2 Implementation Details
Settings on LoNPE Algorithm. As described in Section 3.2.1, to accurately calculate the noise prior of a noisy observation in raw domain, our LoNPE algorithm $\Psi_{\text{LoNPE}}$ is applied on top $\frac{m}{n}=10\%$ of the sampled local patches, which are at size of $\mathcal{O}=16\times 16$ for each raw image. Given a $h\times w$ raw noisy image, we first sample $n=\left\lfloor\frac{h}{16}\right\rfloor\times\left\lfloor\frac{w}{16}\right\rfloor$ local patches, and select the top $m$ smooth patches with lower $\lambda_{\mathcal{S}}$, then calculate its noise prior using Eq.(13).
Settings on LoNPE Network. As described in Section 3.2.2, for practical applications on estimating the noise prior of a sRGB noisy image, we need to train an effective LoNPE network $\Phi_{\text{LoNPE}}$ as illustrated in Fig. 4. Similar to [66], for a single noisy image, we randomly sample 8 local patches of size $32\times 32$ and average the output $(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ as the final predicted noise prior. In the training phase, AdamW optimizer [75] is employed with cosine annealing [76] learning rate from $10^{-3}$ to $10^{-6}$ during 50K mini-batch iterations, on minimizing the objective function in Eq.(15), and the batch size is set to 64.
@@
Settings on Condformer. Following the settings of Restormer in [9], we build our Condformer with groups of Transformer block in the encoder or decoder modules. Yet in the latent module, we stack 8 CondSA blocks with the predicted noise prior to recify the latent space. In the training phase, we adopt the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and set weight decay to $10^{-4}$. Using progressive training strategy proposed by [9], we set the batch size and patch size pairs to [$(64,128^{2})$, $(16,256^{2})$, $(8,384^{2})$, $(4,512^{2})$] at training iterations [0k, 150k, 200k, 250k]. We train our Condformer models for total 300k iterations and the initial learning rate is set to $4\times 10^{-4}$ and gradually reduced to $10^{-6}$ through the cosine annealing. Data augmentation is performed on the training data through horizontal flip and random rotation of 90, 180, and 270.
@@ 4.2 Experiments on Noise Statistics
In this section, we conduct several statistical experiments by analyzing the statistics of synthetic or real noises, to further demonstrate the effectiveness of our LoNPE algorithm on noise prior representation.
@@ 4.2.1 On synthetic noises
As illustrated in Section 3.1.2, the noise prior is beyond and independent of the image prior. From Eq. (5), the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ affects raw noisy observations but is implicit in the sRGB color observations due to unknown ISP operations. To address this, we randomly add noises on the clean sRGB images with Poisson-Gaussian sampling and employ our LoNPE algorithm on the synthesized noisy images to estimate the corresponding noise prior.
As shown in Fig. 6, we set three degree of noise priors $(\sigma_{\text{s}},\sigma_{\text{r}})$ as $(0.05,0.02)$, $(0.10,0.04)$ and $(0.15,0.08)$, representing mild, moderate and severe corruptions, respectively. We sample Poisson-Gaussian noises and add them to 100 images from the Urban100 dataset. By applying LoNPE algorithm, we obtain the estimated noise prior values of each image with different noise level, and find that the mean estimation results are close to the groundtruthes with less discrepancy, particularly for the signal-dependent noise prior $\sigma_{\text{s}}$. However, there exists a nonnegligible discrepancy in the signal-independent noise prior $\sigma_{\text{r}}$ estimation, and the discrepancy increases as the noise level is enlarged. A reason is that, the image context reflects the signal intensity of the target scene illuminance, and could interfere the estimation of the signal-independent $\sigma_{\text{r}}$. This issue is particularly pronounced at the high-frequency regions, e.g., edges and textures. That is why we need to employ the local smoothness criterion $\lambda_{\mathcal{S}}$ to eliminate the interference of image prior.
@@ 4.2.2 On real noises
As described in Section 3.2, the goal of LoNPE is to calculate the noise prior from a single raw noisy observation based on the independence of noise prior and image prior. Thus, the most critical thing for demonstration is the statistical experiment on the correlations of the estimated noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ and the imaging environments.
@@
Method Publisher Noise Params CBSD68 Kodak24 Urban100
Model (M) $\sigma_{\text{s}}$=0 $\sigma_{\text{s}}$=0.1 $\sigma_{\text{s}}$=0.3 $\sigma_{\text{s}}$=0 $\sigma_{\text{s}}$=0.1 $\sigma_{\text{s}}$=0.3 $\sigma_{\text{s}}$=0 $\sigma_{\text{s}}$=0.1 $\sigma_{\text{s}}$=0.3
Noisy - - - 20.12 17.10 13.88 20.02 16.97 13.72 20.27 17.12 13.87
DnCNN [15] TIP’17 G 0.7 31.00 29.10 26.30 31.89 30.12 27.27 30.46 28.70 25.64
Restormer [9] CVPR’22 G 26.1 31.57 28.98 25.40 32.82 30.24 26.62 32.65 29.84 25.68
VDN [16] NeurIPS’19 SV-G 2.0 31.21 29.33 27.22 32.28 30.49 28.33 31.46 29.61 27.05
DRANet [77] PR’24 SV-G 1.6 31.33 29.44 27.30 32.50 30.69 28.44 31.93 30.09 27.65
VIRNet [28] TPAMI’24 SV-G 10.5 31.45 29.52 27.17 32.65 30.76 28.24 32.23 30.24 27.14
FBI-D [78] CVPR’21 P-G 7.5 26.59 25.55 23.30 27.84 26.69 24.11 25.60 24.56 22.32
VBDNet [79] PR’23 P-G 8.3 25.93 25.51 24.41 27.00 26.73 25.68 26.45 25.86 24.57
CLIPDenoising [80] CVPR’24 P-G 34.5 30.52 28.18 25.19 31.31 29.02 25.92 30.04 27.79 24.48
[…]
MambaIR [47] ECCV’24 P-G 23.2 31.41 29.60 27.94 32.65 30.99 29.41 32.42 30.91 29.32
Condformer Ours P-G 27.0 31.59 29.84 28.14 32.81 31.20 29.58 32.64 31.15 29.56
@@
Based on the independence of image prior and noise prior investigated in Fig. 2, the noise prior is quantifiable using our LoNPE algorithm. To demonstrate the effectiveness of our LoNPE algorithm, we conduct a statistical study on the estimated noise priors of SIDD noisy observations with various target scenes and imaging environments. Specifically, according to the camera information of the SIDD-Medium raw-domain training dataset, we mainly analyze the estimated noise prior under various brightness (including “high-light”, “normal-light” and “low-light”), scenes (including scene “001”, “002”, “007” and “008”), ISO levels (ranging from 100 to 3200), and shutter speed (ranging from 1/1600 to 1/60). Subsequently, the statistical results in Fig. 7 can be summarized into several points:
1) Higher ISO level leads to more shot noise. Under same brightness, the shot noise prior $\sigma_{\text{s}}$ is proportional to the ISO levels. Since the ISO settings represent the photo-sensibility of camera sensor, a higher ISO level amplifies the electrical signal generated by the photons hitting the sensor. As shot noise follows Poisson distribution, its relative strength becomes more pronounced as you increase the ISO. The higher the amplification, the more noticeable the noise will be in the final image. This is crucial in low-light conditions to brighten the image. Thus, the slope of variance ($\sigma^{2}_{\text{s}}$) in Eq. (5) would be larger as the ISO increased.
2) Higher shutter speed leads to more read noises. Under same brightness, the read noise prior $\sigma_{\text{r}}$ is proportional to the shutter speed. In camera sensor, especially those designed for high-speed photography, the demand to read out quickly can introduce more noises. Fast readout processes might require higher readout clock speeds and increased electronic noise, contributing to the overall read noise.
3) Independence of noise and image priors. Empirically, the environmental brightness is proportional to the intensity of photons and should influence only shot noise; instead, the ISO and shutter speed indicate the magnification of electrons from photons and would affect the shot noise and read noise, respectively. Besides, the captured signal in a digital image is a sum of both the scene-derived signal and various sensor noises. However, these sensor noises overlay the signal in a way that their behavior and characteristics do not alter based on the attributes of the target scene itself. This phenomenon is well demonstrated by statistical analysis in Fig. 7. We can find that different scenes (with various image priors) have similar noise prior under the same imaging environment, which is strong evidence of demonstrating the independence of noise prior and image prior.
Understanding the independence of noise prior and image prior allows photographers and engineers to develop better noise removal algorithms and improve sensor designs. By treating the scene and sensor noise as separate entities, it becomes easier to process images to enhance the desired signal intensity while minimizing the impact of sensor noise on the final image.

### v2

@@ 3.3 Conditional Denoising Transformer
Due to the strong capability of Transformer-based denoising models (e.g., Uformer 56, Restormer 67, GRL 30 and etc), we design a Conditional denoising Transformer (Condformer) by embedding the noise prior into the self-attention module. Particularly, as the independence of noise prior and image prior, guiding the model to learn from image prior and noise prior separately is the primary principle in designing the Condformer.
@@ 3.3.1 Embedding noise prior in latent space
Existing image denoising networks typically use a global residual connection for noise prediction and an U-shape encoder-decoder structure for feature representation, suggesting that noise is implicitly concealed in the latent space as illustrated in Fig. 5. Meanwhile, image prior such as scene context is theoretically weakest in the latent space. Inspired by this and considering the residual attribute of noise, the latent space code of a noisy image significantly represents noise statistics.
As mentioned earlier in Section 1, an effective conditional denoising model should separately consider the image and noise priors, adhering to the principle of their independence. Therefore, the noise prior should be embedded in the latent space to strengthen noise statistics representation and guide the denoising network to focus more on noisy residues. Specifically as shown in Fig. 5, we firstly extract the latent space code $\mathbf{X}$ of the noisy image using a denoiser encoder, and then construct a feature fusion module to embed the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$.
@@ 3.3.2 Overall pipeline
Based on the principle of embedding the noise prior in the latent space, it is feasible to incorporate the noise prior into any encoder-decoder denoiser. Considering the efficiency in practical applications, we employ the Restormer as our denoiser baseline, replacing its self-attention module by a conditional self-attention module that embeds noise prior in the latent space, while keeping all other modules unchanged. Following Zamir et al 67, given a noisy color observation $\mathbf{I}_{\text{rgb}}\in\mathbb{R}^{3\times h\times w}$, we construct a multi-scale hierarchical denoiser encoder which consists of three channel-wise Transformer blocks to capture cross-covariance across channels, generating the latent space code $\mathbf{X}\in\mathbb{R}^{8c\times\frac{h}{8}\times\frac{w}{8}}$, where $c$ is the number of channels.
@@
${\mathbf{Y}}=\text{CondSA}(\mathbf{X},\mathbf{z})$ (16)
This is then fed into a feed-forward network (FFN) for feature transformation. In particular, to effectively exploit the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ estimated by LoNPE algorithm or network, this prior is firstly encoded into a latent conditional embedding vector $\mathbf{z}\in\mathbb{R}^{1\times c_{z}}$ using a shallow module with fully-connected layers. Indicating single CondSA has a specific embedding vector to adapt the intermediate features. The rectified latent space code equipped with noise prior related embedding, guides the denoiser decoder for specific denoising. In particularly, to preserve fine structural and textural details in the restored images, we use a hierarchical skip-connection strategy 67 that integrates low-level features from the encoder and high-level features from the decoder.
@@ 3.3.3 Conditional Self-Attention Mechanism
where $\mathbf{X}$ and $\mathbf{Y}$ are the input and output features. $\mathbf{Q}\in\mathbb{R}^{c\times hw}$, $\mathbf{K}\in\mathbb{R}^{c\times hw}$ and $\mathbf{V}\in\mathbb{R}^{c\times hw}$ indicate the query, key and value matrix obtained by encoding the input feature $\mathbf{X}\in\mathbb{R}^{c\times h\times w}$ with linear layers $W^{Q}$, $W^{K}$ and $W^{V}$ respectively, each of which stacks a $\times 1$ 1 convolution and a $\times 3$ 3 depth-wise convolution layer. Besides, $\alpha$ is a learnable scaling parameter to control the magnitude of the cross-covariance of $\mathbf{Q}$ and $\mathbf{K}$ before applying a softmax layer.
@@
where $\mathbf{z}\in\mathbb{R}^{1\times 2c_{z}}$ represents the conditional embedding vector from the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ by repeating $k$ times in channel dimension.
The conditional attention should effectively represent the correlation between-in the intermediate image features, and the latent correlation between the intermediate image features and noise prior. Therefore, as mentioned in Section 3.1.1, a feature fusion module is essential for capturing the relationship between the noise prior and latent code. Intuitively, the query tensor $\mathbf{Q}$ and the key tensor $\mathbf{K}$ indicates the information for feature retrieval 52; instead, the value tensor $\mathbf{V}$ represent the property of the input feature. Therefore, it is reasonable to embed noise prior into the query/key tensors and generate the conditional counterparts $\mathbf{Q}^{\prime}$ and $\mathbf{K}^{\prime}$, so we have
@@
where $\Phi_{\text{LFM}}$ represents a linear fusion module (LFM) layer for feature fusion. In particular, since noise is distributed across both spatial and channel dimensions of intermediate image features, the LFM fuses the query/key tensors and the conditional embedding vector in both dimensions by concatenating them along the channel axis and applying a $\times 1$ 1 convolution and a $\times 3$ 3 depth-wise convolution layer. This design enables effective and localized integration of noise prior into the feature representations.
@@ 4 Experiments
We describe the experimental setup and then evaluate the performances of our LoNPE and Condformer on noise statistics and blind image denoising. Finally, we perform ablation studies to demonstrate the effectiveness of our methods.
@@ 4.1.1 Datasets and Metrics
Synthetic Datasets. Following 73, we adopt several sRGB image datasets for training the LoNPE and Condformer networks, including the DIV2K and Flicker2K dataset 2 with 3650 high-quality 2K images, the Berkeley segmentation dataset (BSD) 41 with 400 images, Waterloo Exploration Database (WED) dataset 37 with 4744 images. Considering both signal-dependent and signal-independent noises, we randomly add noises on the clean sRGB image with Poisson-Gaussian noise model. The noise level are set to $\sigma_{\text{s}}\sim\mathbb{U}(0,0.3)$ and $\sigma_{\text{r}}\sim\mathbb{U}(0,50/255)$, both of which indicate the groundtruth noise prior to demonstrate the effectiveness of our LoNPE algorithm. Especially, the Poisson-Gaussian noise will be degraded to additive Gaussian white noise (AWGN) when $\sigma_{\text{s}}=0$. To evaluate the performance, we apply our methods on several benchmarks, including CBSD68 41, Kodak24 19 and Urban100 25.
Real Datasets. Consistent with previous real image denoising work 67, we adopt the SIDD-Medium dataset 1 for real noise statistics and real image denoising tasks, which contains 320 noisy-clean image pairs in both raw and sRGB domains. Particularly, we first apply our LoNPE algorithm on 320 noisy raw images to estimate their corresponding noise priors, and then utilize them to help training the LoNPE and Condformer networks on the 320 sRGB noisy-clean image pairs. For validation, 1024 pairs of noisy-clean sRGB image patches from SIDD validation dataset 1 are adopted. Besides, evaluation is also conducted on 1280 noisy $\times 256$ 256 patches from the SIDD benchmark dataset 1 and 50 noisy $\times 512$ 512 images from the DND benchmark dataset 46.
Evaluation Metrics. Two commonly-used image quality assessment criteria are adopted to evaluate the performances: Peak Signal-to-Noise Ratio (PSNR) and Structural SIMilarity (SSIM) 55, and are calculated in the sRGB domain. Note that, since the clean groundtruthes of SIDD and DND benchmark datasets are unavailable, we calculate these metrics on only SIDD validation dataset and obtain the metrics of them from the online servers [^1] SIDD: https://abdokamel.github.io/sidd/;
@@
Table 1: Comparative study of noise prior estimation on the Urban100 dataset. Random noises are sampled based on the given noise prior parameters $(\sigma_{\text{s}},\sigma_{\text{r}})$ and added to each clean image. Note that the top three methods are executed on a CPU platform, while the bottom three methods are executed on a GPU platform.
Methods Noise prior parameter estimation result of $(\sigma_{\text{s}},\sigma_{\text{r}})$ Time (s)
Gaussian Poisson Poisson-Gaussian
(0.000, 0.050) (0.100, 0.000) (0.050, 0.020) (0.100, 0.050) (0.200, 0.100)
Mäkitalo and Foi 39 (0.441, 0.044) (0.111, 0.006) (0.186, 0.018) (0.446, 0.041) (0.866, 0.089) 56.32
Pimpalkhute et al 45 ( - , 0.039) ( - ,  - ) ( - , 0.021) ( - , 0.039) ( - , 0.075) 0.08
LoNPE Algorithm (Ours) (0.053, 0.066) (0.104, 0.009) (0.062, 0.029) (0.117, 0.057) (0.233, 0.091) 0.17
[…]
LoNPE Network (Ours) (0.022, 0.043) (0.088, 0.015) (0.046, 0.020) (0.092, 0.045) (0.182, 0.075) 0.01
Figure 6: Statistical study on noise prior estimation performance of 39 and our LoNPE algorithm with synthetic Poisson-Gaussian noises. By applying our LoNPE algorithm, we can effectively calculate a relative accurate noise prior from a single noisy observation.
@@ 4.1.2 Implementation Details
Settings on LoNPE Algorithm. As described in Section 3.2.1, to accurately calculate the noise prior of a noisy observation in raw domain, our LoNPE algorithm $\Psi_{\text{LoNPE}}$ is applied on top $\frac{m}{n}=10\%$ of the sampled local patches, which are at size of $\mathcal{O}=16\times 16$ for each raw image. Given a $h\times w$ raw noisy image and set the sampling stride to be $k=4$, we first sample $n=\left\lfloor\frac{h}{k}\right\rfloor\times\left\lfloor\frac{w}{k}\right\rfloor$ local patches, and select the top $m$ smooth patches with lower $\lambda_{\mathcal{S}}$, then calculate its noise prior using Eq.(13).
Settings on LoNPE Network. As described in Section 3.2.2, for practical applications on estimating the noise prior of a sRGB noisy image, we need to train an effective LoNPE network $\Phi_{\text{LoNPE}}$ as illustrated in Fig. 4. Similar to 26, for a single noisy image, we randomly sample 8 local patches of size $\times 32$ 32 and average the output $(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ as the final predicted noise prior. In the training phase, AdamW optimizer 36 is employed with cosine annealing 35 learning rate from $10^{-3}$ to $10^{-6}$ during 50K mini-batch iterations, on minimizing the objective function in Eq.(15), and the batch size is set to 64.
[…]
@@ 4.2 Experiments on Noise Statistics
In this section, we conduct several statistical experiments to verify the effectiveness of our LoNPE algorithm and network on noise prior representation. In particular, we first conduct noise prior estimation experiment on synthetic Poisson-Gaussian noises quantitatively, and further analyze the statistics of real noise.
@@ 4.2.1 On synthetic noises
As illustrated in Section 3.1.2, the noise prior is beyond and independent of the image prior. From Eq. (5), the noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ affects raw noisy observations but is implicit in the sRGB color observations due to unknown ISP operations. To address this limitation, we randomly add noises to clean sRGB images with Poisson-Gaussian sampling and employ our LoNPE algorithm on the synthesized noisy images to estimate the corresponding noise prior.
Quantitatively, we perform a comparative study on noise prior parameter estimation, as reported in Table 1. By setting various noise prior parameters, Gaussian, Poisson and the complicated Poisson-Gaussian noise types are considered. For the comparison of synthetic noise prior estimation, several noise parameter estimation methods are evaluated, including the traditional method for Gaussian 45 or Poisson-Gaussian noises 39, and the CNN-based Poisson-Gaussian noise parameter estimators in FBI-Denoiser 7 with fixed and mixed noise levels. Particularly for Poisson-Gaussian noises, it is observed that Mäkitalo and Foi 39 fails when signal-independent noise is more severe than signal-dependent noise, and commonly requires significant computation time to search for intersection of the unitary variance contours. In contrast, our LoNPE algorithm effectively handles more general scenarios, including Gaussian, Poisson and Gaussian-Poisson noises, with a speedup of up to ×300$\times$ 300× 300. Additionally, by leveraging GPU acceleration, our LoNPE network increases speed significantly with negligible performance degradation then LoNPE algorithm, and achieves superior generalization across all scenarios compared to the noise parameter estimator in FBI-Denoiser 7.
To further demonstrate the stability of our LoNPE algorithm, we visualize the statistical histogram of the estimated noise priors. As shown in Fig. 6, we set three levels of noise prior parameter $(\sigma_{\text{s}},\sigma_{\text{r}})$ as (0.05,0.02)(0.05,0.02)( 0.05 , 0.02 ), (0.10,0.05)(0.10,0.05)( 0.10 , 0.05 ) and (0.2,0.1)(0.2,0.1)( 0.2 , 0.1 ), which is consistent with the setting in Table 1, representing mild, moderate and severe corruptions, respectively. By applying LoNPE algorithm, we obtain the estimated noise prior parameters of each image with different noise level, and find that the mean values of estimation results are closer to the groundtruthes than the compared method Mäkitalo and Foi 39, particularly for the signal-dependent noise prior $\sigma_{\text{s}}$. Nevertheless, there exists a nonnegligible discrepancy in the signal-independent noise prior $\sigma_{\text{r}}$ estimation, and the discrepancy increases as the noise level is enlarged. A reason is that, the image context reflects the signal intensity of the target scene illuminance, and could interfere the estimation of the signal-independent $\sigma_{\text{r}}$. This issue is particularly pronounced at the high-frequency regions, e.g., edges and textures. That is why we need to employ the local smoothness criterion $\lambda_{\mathcal{S}}$ to eliminate the interference of image prior.
@@ 4.2.2 On real noises
Furthermore, as described in Section 3.2, the goal of LoNPE is to calculate the noise prior from a single raw noisy observation based on the independence of noise prior and image prior. Thus, the most critical effectiveness demonstration for real scenes is a statistical experiment on the correlations of the estimated noise prior $(\sigma_{\text{s}},\sigma_{\text{r}})$ and the camera sensor imaging environments.
Based on the independence of image prior and noise prior investigated in Fig. 2 and the statistical study on synthetic Poisson-Gaussian noise prior estimation in Section 4.2.1, the noise prior is quantifiable using our LoNPE algorithm. Thus, we conduct a statistical study on the estimated noise priors of SIDD noisy observations with various target scenes and imaging environments. Specifically, according to the camera information of the SIDD-Medium raw-domain training dataset 1, we mainly analyze the estimated noise prior under various brightness (including “high-light”, “normal-light” and “low-light”), scenes (including scene “001”, “002”, “007” and “008”), ISO levels (ranging from 100 to 3200), and shutter speed (ranging from 1/1600 to 1/60). Subsequently, the statistical results in Fig. 7 can be summarized into several points:
1) Higher ISO level leads to more shot noise. Shot noise prior $\sigma_{\text{s}}$ grows with ISO levels, as higher ISO amplifies the signal generated by photons on the camera sensor. Since shot noise follows a Poisson distribution, amplification makes the noise more pronounced, especially in low-light conditions, affecting image quality. Thus, the slope of variance ($\sigma^{2}_{\text{s}}$) in Eq. (5) would be larger as the ISO increased.
2) Higher shutter speed leads to more read noises. Under same brightness, the read noise prior $\sigma_{\text{r}}$ scales with shutter speed. Faster shutter speeds, particularly in high-speed photography, require rapid sensor readout, which introduces additional electronic noise, leading to higher read noise.
3) Independence of noise and image priors. Brightness influences shot noise via photon intensity, while ISO and shutter speed affect shot and read noise, respectively, without altering the scene-derived signal. Statistical analysis confirms that noise characteristics remain consistent across different scenes under the same imaging conditions, proving the independence of noise priors from image priors.
@@
Method Noise Params CBSD68 Kodak24 Urban100
Model (M) $\sigma_{\text{s}}$=0 $\sigma_{\text{s}}$=0.15 $\sigma_{\text{s}}$=0.3 $\sigma_{\text{s}}$=0 $\sigma_{\text{s}}$=0.15 $\sigma_{\text{s}}$=0.3 $\sigma_{\text{s}}$=0 $\sigma_{\text{s}}$=0.15 $\sigma_{\text{s}}$=0.3
Noisy - - 20.12 17.10 13.88 20.02 16.97 13.72 20.27 17.12 13.87
DnCNN 71 G 0.7 31.00 29.10 26.30 31.89 30.12 27.27 30.46 28.70 25.64
Restormer 67 G 26.1 31.57 28.98 25.40 32.82 30.24 26.62 32.65 29.84 25.68
VDN 63 SV-G 2.0 31.21 29.33 27.22 32.28 30.49 28.33 31.46 29.61 27.05
DRANet 59 SV-G 1.6 31.33 29.44 27.30 32.50 30.69 28.44 31.93 30.09 27.65
VIRNet 64 SV-G 10.5 31.45 29.52 27.17 32.65 30.76 28.24 32.23 30.24 27.14
FBI-Denoiser 7 P-G 7.5 26.59 25.55 23.30 27.84 26.69 24.11 25.60 24.56 22.32
VBDNet 31 P-G 8.3 25.93 25.51 24.41 27.00 26.73 25.68 26.45 25.86 24.57
CLIPDenoising 13 P-G 34.5 30.52 28.18 25.19 31.31 29.02 25.92 30.04 27.79 24.48
Restormer 67 P-G 26.1 31.48 29.66 28.01 32.70 31.02 29.44 32.46 30.95 29.37
MambaIR 21 P-G 23.2 31.41 29.60 27.94 32.65 30.99 29.41 32.42 30.91 29.32
Condformer (Ours) P-G 27.0 31.59 29.84 28.14 32.81 31.20 29.58 32.64 31.15 29.56
@@
Consequently, understanding the independence of noise prior and image prior allows photographers and engineers to develop better noise removal algorithms and improve sensor designs. By treating the scene and sensor noise as separate entities, it becomes easier to process images to enhance the desired signal intensity while minimizing the impact of sensor noise on the final image.